OpenStudy (anonymous):

Please help me Find the following limits: 1. limit (as x approaches infinity) :(sqrt(2(x^2)+x)-sqrt(2)x) 2. lim (as x approaches infinity) : 10^9(sqrt(2(x^2)+x) - sqrt(2)x) 3. lim(as x approaches neg infinity) : (sqrt(2(x^2) +3))/x

OpenStudy (rsadhvika):

have you tried rationalizing the numerator for #1 ?

OpenStudy (anonymous):

yes and then i got super lost

OpenStudy (rsadhvika):

\[\begin{align} \lim\limits_{x\to\infty} \sqrt{2x^2+x} - \sqrt{2}x \end{align}\]

OpenStudy (rsadhvika):

is that the first limit ?

OpenStudy (anonymous):

yes

OpenStudy (rsadhvika):

\[\begin{align} \lim\limits_{x\to\infty} \sqrt{2x^2+x} - \sqrt{2}x & = \lim\limits_{x\to\infty} \sqrt{2x^2+x} - \sqrt{2}x\times\dfrac{ \sqrt{2x^2+x} + \sqrt{2}x}{\sqrt{2x^2+x} + \sqrt{2}x}\\~\\ & = \lim\limits_{x\to\infty} \dfrac{\left(\sqrt{2x^2+x}\right)^2 - \left(\sqrt{2}x\right)^2}{\sqrt{2x^2+x} + \sqrt{2}x}\\~\\ & = \lim\limits_{x\to\infty} \dfrac{x}{\sqrt{2x^2+x} + \sqrt{2}x}\\~\\ \end{align}\]

OpenStudy (rsadhvika):

next divide top and bottom by x

OpenStudy (anonymous):

so 1/((sqrt(2x^2+x) /x) +sqrt (2)?

OpenStudy (rsadhvika):

\[\begin{align} \lim\limits_{x\to\infty} \sqrt{2x^2+x} - \sqrt{2}x & = \lim\limits_{x\to\infty} \sqrt{2x^2+x} - \sqrt{2}x\times\dfrac{ \sqrt{2x^2+x} + \sqrt{2}x}{\sqrt{2x^2+x} + \sqrt{2}x}\\~\\ & = \lim\limits_{x\to\infty} \dfrac{\left(\sqrt{2x^2+x}\right)^2 - \left(\sqrt{2}x\right)^2}{\sqrt{2x^2+x} + \sqrt{2}x}\\~\\ & = \lim\limits_{x\to\infty} \dfrac{x}{\sqrt{2x^2+x} + \sqrt{2}x}\\~\\ & = \lim\limits_{x\to\infty} \dfrac{x \color{red}{/x}}{\sqrt{2x^2+x} \color{red}{/x} + \sqrt{2}x \color{red}{/x}}\\~\\ \end{align}\]

OpenStudy (rsadhvika):

when u push that x inside radical, it becomes x^2

OpenStudy (anonymous):

so sqrt (2x^4+x)?

OpenStudy (rsadhvika):

\[\begin{align} \lim\limits_{x\to\infty} \sqrt{2x^2+x} - \sqrt{2}x & = \lim\limits_{x\to\infty} \sqrt{2x^2+x} - \sqrt{2}x\times\dfrac{ \sqrt{2x^2+x} + \sqrt{2}x}{\sqrt{2x^2+x} + \sqrt{2}x}\\~\\ & = \lim\limits_{x\to\infty} \dfrac{\left(\sqrt{2x^2+x}\right)^2 - \left(\sqrt{2}x\right)^2}{\sqrt{2x^2+x} + \sqrt{2}x}\\~\\ & = \lim\limits_{x\to\infty} \dfrac{x}{\sqrt{2x^2+x} + \sqrt{2}x}\\~\\ & = \lim\limits_{x\to\infty} \dfrac{x \color{red}{/x}}{\sqrt{2x^2+x} \color{red}{/x} + \sqrt{2}x \color{red}{/x}}\\~\\ & = \lim\limits_{x\to\infty} \dfrac{1}{\sqrt{2x^2\color{red}{/x^2}+x\color{red}{/x^2}} + \sqrt{2}} \end{align}\]

OpenStudy (anonymous):

oh, ok,

OpenStudy (anonymous):

so 1/(sqrt(2+1/x) + sqrt (2)?

OpenStudy (rsadhvika):

\[\begin{align} \lim\limits_{x\to\infty} \sqrt{2x^2+x} - \sqrt{2}x & = \lim\limits_{x\to\infty} \sqrt{2x^2+x} - \sqrt{2}x\times\dfrac{ \sqrt{2x^2+x} + \sqrt{2}x}{\sqrt{2x^2+x} + \sqrt{2}x}\\~\\ & = \lim\limits_{x\to\infty} \dfrac{\left(\sqrt{2x^2+x}\right)^2 - \left(\sqrt{2}x\right)^2}{\sqrt{2x^2+x} + \sqrt{2}x}\\~\\ & = \lim\limits_{x\to\infty} \dfrac{x}{\sqrt{2x^2+x} + \sqrt{2}x}\\~\\ & = \lim\limits_{x\to\infty} \dfrac{x \color{red}{/x}}{\sqrt{2x^2+x} \color{red}{/x} + \sqrt{2}x \color{red}{/x}}\\~\\ & = \lim\limits_{x\to\infty} \dfrac{1}{\sqrt{2x^2\color{red}{/x^2}+x\color{red}{/x^2}} + \sqrt{2}} \\~\\ & = \lim\limits_{x\to\infty} \dfrac{1}{\sqrt{2+1/\color{red}{x}} + \sqrt{2}} \end{align}\]

OpenStudy (rsadhvika):

yes take the limit now

OpenStudy (rsadhvika):

as x -> infinity 1/x -> 0

OpenStudy (rsadhvika):

\[\begin{align} \lim\limits_{x\to\infty} \sqrt{2x^2+x} - \sqrt{2}x & = \lim\limits_{x\to\infty} \sqrt{2x^2+x} - \sqrt{2}x\times\dfrac{ \sqrt{2x^2+x} + \sqrt{2}x}{\sqrt{2x^2+x} + \sqrt{2}x}\\~\\ & = \lim\limits_{x\to\infty} \dfrac{\left(\sqrt{2x^2+x}\right)^2 - \left(\sqrt{2}x\right)^2}{\sqrt{2x^2+x} + \sqrt{2}x}\\~\\ & = \lim\limits_{x\to\infty} \dfrac{x}{\sqrt{2x^2+x} + \sqrt{2}x}\\~\\ & = \lim\limits_{x\to\infty} \dfrac{x \color{red}{/x}}{\sqrt{2x^2+x} \color{red}{/x} + \sqrt{2}x \color{red}{/x}}\\~\\ & = \lim\limits_{x\to\infty} \dfrac{1}{\sqrt{2x^2\color{red}{/x^2}+x\color{red}{/x^2}} + \sqrt{2}} \\~\\ & = \lim\limits_{x\to\infty} \dfrac{1}{\sqrt{2+1/\color{red}{x}} + \sqrt{2}}\\~\\ & = \dfrac{1}{\sqrt{2+\color{red}{0}} + \sqrt{2}}\\~\\ \end{align}\]

OpenStudy (rsadhvika):

simplify

OpenStudy (anonymous):

so as x approaches infinity the limit would be going to 1/sqrt(2) right? Thanks so much!!

OpenStudy (rsadhvika):

try again

OpenStudy (rsadhvika):

in the bottom there are two radicals right

OpenStudy (anonymous):

oh wait 1/(2sqrt(2)

OpenStudy (rsadhvika):

yes!

OpenStudy (anonymous):

Thanks that really helped me understand what Im trying to get to

OpenStudy (rsadhvika):

for #2, simply multiply this by 10^9

OpenStudy (anonymous):

so on the second one, when you add the 10^9 out in front wouldnt it just be 10^9(1/(ssqrt2) or am I not thinking about that correctly?

OpenStudy (rsadhvika):

you're right, if the limit exists, : \[\large \lim\limits_{x\to } Cf(x) =C\lim\limits_{x\to } f(x) \]

OpenStudy (anonymous):

great, and on the last one can't you just think that the top is approaching infinity and bottom is approaching neg infinity so the whole thing is approaching -infinity? Or am I way off?

OpenStudy (rsadhvika):

\[ \lim\limits_{x\to\infty}10^9 \left( \sqrt{2x^2+x} - \sqrt{2}x\right) = 10^9 \lim\limits_{x\to\infty} \left( \sqrt{2x^2+x} - \sqrt{2}x\right) = 10^9 \left(\dfrac{1}{2\sqrt{2}}\right)\]

OpenStudy (rsadhvika):

you have good intuition, but its tricky.. u need to work it using limit properties

OpenStudy (rsadhvika):

you will be in problem if the denominator approaches inifnity at faster rate compared to numerator

OpenStudy (rsadhvika):

then, the aggressive denominator pulls the entire fraction to 0 and the limit also will be 0

OpenStudy (rsadhvika):

intuition is good, but it wont work always with limits

OpenStudy (rsadhvika):

lets divide top and bottom by x and see what we get

OpenStudy (anonymous):

Thats what i was just working on :) I got sqrt (2+ 3/x)

OpenStudy (rsadhvika):

no wait, i was wrong.. lets first make a substitution

OpenStudy (anonymous):

doesn't 3/x approach 0? and we are left with 2 or am i looking at that wrong?

OpenStudy (rsadhvika):

lets turn negative infinity to positive infinity

OpenStudy (anonymous):

ok

OpenStudy (rsadhvika):

substitute x = -y as x-> -infinity, y -> infinity right ?

OpenStudy (rsadhvika):

\[\begin{align} \lim\limits_{x\to-\infty} \dfrac{\sqrt{2x^2+3}}{x} &= \lim\limits_{y\to\infty} \dfrac{\sqrt{2(-y)^2+3}}{(-y)} \end{align}\]

OpenStudy (rsadhvika):

are you happy with above ? (you will see why it is needed shortly)

OpenStudy (rsadhvika):

\[\begin{align} \lim\limits_{x\to-\infty} \dfrac{\sqrt{2x^2+3}}{x} &= \lim\limits_{y\to\infty} \dfrac{\sqrt{2(-y)^2+3}}{(-y)} \\~\\ &= \lim\limits_{y\to\infty} \dfrac{\sqrt{2y^2+3}}{-y} \\~\\ \end{align}\]

OpenStudy (anonymous):

So you can change the infinity by taking the opposite of the variable?

OpenStudy (rsadhvika):

i have just substituted x by -y

OpenStudy (rsadhvika):

everywhere

OpenStudy (rsadhvika):

since x is going to -infinity, y has to go to +infinity right ?

OpenStudy (rsadhvika):

\[\begin{align} \lim\limits_{x\to-\infty} \dfrac{\sqrt{2x^2+3}}{x} &= \lim\limits_{y\to\infty} \dfrac{\sqrt{2(-y)^2+3}}{(-y)} \\~\\ &= \lim\limits_{y\to\infty} \dfrac{\sqrt{2y^2+3}}{-y} \\~\\ &= \lim\limits_{y\to\infty} -\sqrt{2+3/y} \\~\\ \end{align}\]

OpenStudy (rsadhvika):

take the limit now

OpenStudy (rsadhvika):

as y -> infinity, 1/y -> 0 so u get : \[\begin{align} \lim\limits_{x\to-\infty} \dfrac{\sqrt{2x^2+3}}{x} &= \lim\limits_{y\to\infty} \dfrac{\sqrt{2(-y)^2+3}}{(-y)} \\~\\ &= \lim\limits_{y\to\infty} \dfrac{\sqrt{2y^2+3}}{-y} \\~\\ &= \lim\limits_{y\to\infty} -\sqrt{2+3/y} \\~\\ &= -\sqrt{2+0} \\~\\ \end{align}\]

OpenStudy (rsadhvika):

simplify

OpenStudy (anonymous):

so it is -sqrt 2?

OpenStudy (rsadhvika):

yup!

OpenStudy (rsadhvika):

u may use wolfram to double check ur answers in future http://www.wolframalpha.com/input/?i=+lim%28as+x+approaches+neg+infinity%29+++++%28sqrt%282%28x%5E2%29+%2B3%29%29%2Fx

OpenStudy (anonymous):

Thanks, so we added the negative because it was to negative infinity originally?

OpenStudy (rsadhvika):

yes you can use that trick for evaluating limits at neg infinity

OpenStudy (rsadhvika):

convert it to pos inifnity by using that substitution x = -y and work it

OpenStudy (anonymous):

thank you!!!!!